OpenStudy (anonymous):

Rearrange this into quadratic formula step by step 5.333 = (0.6 + 2p)² / (0.135 - p)

OpenStudy (anonymous):

The best thing to start with would be to multiply (0.135-p) with (5.333) so that all p values won't be in the denominator of a fraction. So \[5.333(0.135-p)=(0.6+2p)\]

OpenStudy (anonymous):

Then use distribution and foil out the right hand side of the equation \[0.719955-5.333p=(0.6+2p)(0.6+2p)\]

OpenStudy (anonymous):

0.719955-5.333p=0.36+1.2p+1.2p+4p^2

OpenStudy (anonymous):

Combine like terms \[4p ^{2}+2.4p+0.36+5.333p-0.719955\] \[4p ^{2}+7.733p-0.359955\]

OpenStudy (anonymous):

this is the quadratic expression. Set the expression = 0 to get the quadratic equation

OpenStudy (anonymous):

Thank you very much! This helped a lot!